# Galactic Trajectories of Interstellar Objects 1I/’Oumuamua, 2I/Borisov, and 3I/Atlas

### v4

@@ 4 Conclusions
We have integrated the Galactic trajectories of known interstellar objects back in time. The vertical extent of the trajectory of 1I/’Oumuamua suggests that it originated near the disk mid-plane of the thin disk, where the ages of most stars are below 1–2 Gyr. In contrast, the comet 2I/Borisov’s vertical excursion is similar to that of the Sun, suggesting a similar age. Given its high vertical excursions, the putative source star of 3I/ATLAS is likely older than that of 1I/’Oumuamua or 2I/Borisov.
@@
Our constraints on the ages of the various objects represent upper limits because the velocity dispersion of the interstellar objects includes both the velocity dispersion of their parent stars and the dispersion in their characteristic ejection speed away from their birth system.

### v5

@@ 4 Conclusions
We have developed a comprehensive framework for constraining the origins of interstellar objects through galactic orbital analysis and Bayesian age inference. Our work advances beyond previous single-orbit studies by implementing Monte Carlo uncertainty propagation and rigorous statistical methodology that addresses observational limitations and provides quantitative age estimates with proper confidence intervals.
We performed ensemble calculations using 10,000 orbital realizations per object to properly account for measurement uncertainties in both interstellar object velocities and Solar motion parameters. This approach yielded robust statistical distributions of maximum vertical excursions: $z_{\text{max}}=0.016\pm 0.002$ kpc for 1I/’Oumuamua, $0.121\pm 0.010$ kpc for 2I/Borisov, and $0.480\pm 0.020$ kpc for 3I/ATLAS. The relatively small fractional uncertainties (5-15%) validate the robustness of our orbital calculations while providing essential error bars for subsequent statistical analysis.
We implemented a physically motivated statistical framework that combines the orbital $z_{\text{max}}$ measurements with theoretical models of stellar kinematics. The methodology employs a Rayleigh-like likelihood function derived from first principles, incorporating age-dependent velocity dispersions that capture the observed thin-thick disk transition in galactic stellar populations. Our analysis yields quantitative age estimates: 1I/’Oumuamua originated from a young stellar system (1.0 Gyr, 68% CI: 0.1-4.1 Gyr), 2I/Borisov from an intermediate-age population (3.8 Gyr, 68% CI: 1.8-5.9 Gyr), and 3I/ATLAS from an ancient thick-disk source (9.6 Gyr, 68% CI: 7.8-10.3 Gyr).
The results demonstrate clear age discrimination where smaller vertical excursions correspond systematically to younger stellar origins. 1I/’Oumuamua’s extremely confined orbit indicates ejection from a recently formed stellar system in the thin galactic disk with minimal dynamical heating. 2I/Borisov’s moderate excursions suggest origin from a Solar-age thin disk population with intermediate levels of gravitational scattering. 3I/ATLAS’s large vertical range requires high velocity dispersions characteristic of the ancient thick disk component formed during early violent galactic evolution.
Our dual approach combining traditional scale-height demographics with Bayesian kinematics provides complementary perspectives on interstellar object origins. The Bayesian framework addresses key limitations of demographic methods by incorporating theoretical understanding of stellar velocity evolution and providing rigorous uncertainty quantification essential for robust scientific inference.
@@
Our constraints on the ages of the various objects represent upper limits because the velocity dispersion of the interstellar objects includes both the velocity dispersion of their parent stars and the dispersion in their characteristic ejection speed away from their birth system. However, the Bayesian framework properly accounts for these systematic effects through the likelihood function formulation and provides realistic credible intervals that reflect both observational and theoretical uncertainties.
The diversity of inferred ages across the three objects suggests multiple formation pathways and ejection mechanisms operating throughout galactic history. Young objects like 1I/’Oumuamua may originate from active star formation regions, while ancient objects like 3I/ATLAS likely represent products of early planetary system evolution in the thick disk. This demographic diversity provides valuable constraints on the temporal evolution of planetary system architectures and ejection processes over cosmic time scales.